OpenStudy (anonymous):

Rewrite sin^4(x) so that it involves only the first power of cosine.

OpenStudy (anonymous):

\[\large\begin{align*} \sin^4x&=\left(\sin^2x\right)^2\\ &=\left(\frac{1}{2}-\frac{1}{2}\cos2x\right)^2\\ &=\frac{1}{4}-\frac{1}{2}\cos2x+\frac{1}{4}\cos^22x\\ &=\frac{1}{4}-\frac{1}{2}\cos2x+\frac{1}{4}\left(\frac{1}{2}+\frac{1}{2}\cos4x\right)\\ &=\frac{1}{4}-\frac{1}{2}\cos2x+\frac{1}{8}+\frac{1}{8}\cos4x\\ &=\frac{3}{8}-\frac{1}{2}\cos2x+\frac{1}{8}\cos4x\\ \end{align*}\]

OpenStudy (anonymous):

Hm, could I have some explanation for the steps you took to get to the solution?

OpenStudy (cwrw238):

that was clever in second line he used the following trig identity (this is how its usually given in the books) cos 2x = 1 - 2sin^2 x rearranging we get sin^2 x = , = 1/2( 1 - cos 2x) or 1/2 - 1/2 cos 2x

OpenStudy (cwrw238):

then after expanding he still had a term 1/4 cos^2 2x (line 3)

OpenStudy (cwrw238):

OK?

OpenStudy (anonymous):

Yeah, I get it now. :D

OpenStudy (cwrw238):

then in line 4 he used the identity cos 2x = 2 cos^2 x - 1 and rearranged to make cos^2 x the subject